# Floquet-engineered Emergent Massive Nambu–Goldstone Modes

### v3

@@ Abstract
SM 1 Continuous Driving Protocol
@@
$Q^{(0)}\propto K_{0},\ Q^{(1)}\propto T\sum_{l=1}^{\infty}\frac{1}{l}[K_{l},K_{-l}],$ (S.2)
@@
where $K_{0}$ preserves a symmetry group $G_{2}$, $K_{1}$ and $K_{1}^{\prime}$ are two non-commuting generators of $G_{2}$, and $K_{2}$ further breaks all symmetries. Then the effective Hamiltonian reduces to $Q^{(0)}\propto K_{0},\ Q^{(1)}\propto iT[K_{1},K_{1}^{\prime}]$ and higher-order terms break all symmetries. Hence, this protocol can effectively realize similar mNGs as in the main text but with continuous drives.
@@
$\dot{S}_{j}^{z}=$ $0$
This can be solved analytically by performing a Fourier transform, $S_{j}^{\alpha}{=}\sum_{q}S_{q}^{\alpha}e^{i(qa)j}$, $\alpha{\in}\{x,y,z\}$, with $q$ denoting the quasi-momentum and $a$ being the lattice constant. Finally we get the spin-wave spectrum, Eq. (5)
@@ 2 Oscillating Amplitudes of Quasi-conserved Quantities in Classical Spin Systems
Drive-induced terms $Q^{(2)}$ of order $\mathcal{O}(T^{2})$ break the $O(2)$ symmetry along the $z$-axis. Therefore, in the linearized regime where spins are mostly polarized in the $z$ direction, $Q^{(2)}$ effectively generate weak fields in the $(x,y)$-plane. One can approximately describe this situation via the Hamiltonian
$H=-J_{2}\sum_{j}\bm{S}_{j}\cdot\bm{S}_{j+1}-h(h_{x}\sum_{j}S_{j}^{x}+h_{y}\sum_{j}S_{j}^{y}+h_{z}\sum_{j}S_{j}^{z}),$ (S.7)
@@
$\bm{\dot{S}_{q}}=[2J_{2}(1-\cos{qa})+h]\begin{pmatrix}0&h_{z}&-h_{y}\\ -h_{z}&0&h_{x}\\ h_{y}&-h_{x}&0\\ \end{pmatrix}\bm{S_{q}}\equiv A_{q}\bm{S_{q}}.$ (S.9)
@@
suggesting an oscillatory dynamics with frequency $\omega=2J_{2}(1-\cos{qa})+h$. It is important to notice that, although the field strength in the $(x,y)$-plane is much smaller than the $z$-field, the oscillations in three directions couple together and share the same frequency. However, the oscillating amplitudes are different, if the initial states are almost polarized in the $z$-direction, $S_{q}^{x/y}\ll 1$. More precisely, as long as $S_{q}^{x/y}$ is of order $\mathcal{O}(T)$ or even smaller, the third column of the matrix in Eq. (S.11) will dominate the dynamics, and the oscillating amplitudes scale as $A_{x/y}\sim T$ and $A_{z}\sim T^{2}$.
@@ 3 Asymmetric Lineshape and Side Peaks
The DSF has an asymmetric lineshape, as shown in Fig. 4(a). Also, several side peaks can be clearly resolved, one intriguing and rare non-equilibrium phenomenon in systems without external driving. We conjecture that the side peaks originate from $Q^{(3)}$ of order $\mathcal{O}(T^{3})$, a high-order and weak process. Also, it can be slightly enhanced by Trotter decomposition being used to approximate the Heisenberg Hamiltonian evolution. To demonstrate this, we compare the numerical simulation generated by $Q_{\text{eff}}$ and different order Trotter decomposition.
Figure S1: (a) The splitting between the main peak and the highest side peak $\Delta\omega$ exhibits a power-law scaling as $\Delta\omega\sim T^{\alpha}$, $\alpha\approx 3$, suggesting that these peaks may be generated by $Q^{(3)}$ of order $\mathcal{O}(T^{3})$. (b) Fourier peaks obtained by various numerical methods. We use $T{=}1.2$ and $J_{x}{=}1$. The lineshape becomes asymmetric once $Q^{(2)}$ is included. Fine structures in the lineshape appear in Floquet driven systems. 3rd-order Trotterization leads to Fourier spectrum that is almost identical to those obtained without the Trotterization. 2nd order Trotter slightly enhances the side peaks, although the overall linewidth remains approximately unchanged. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $\theta_{j}{=}0.25\pi$ and $\phi_{j}$ randomly sampled within $(-\pi/2,\pi/2)$. $100$ ensemble averages are performed to the system with its size $N=1000$.
We extract the splitting between the main peak and the highest side peak $\Delta\omega$ and plot it in Fig. S1(a) for different driving period. $\Delta\omega$ exhibits a power-law dependence on the driving period $T$ with the exponent $\alpha\approx 3$.
We try to reproduce the interesting lineshape by several different numerical methods:
@@
2. Floquet driving the system and approximating the Heisenberg Hamiltonian $H_{2}$ using Trotter decomposition of different orders. Details of this method will be explained later in Sec. SM 4.
@@
The results are shown in Fig. S1(b). For $Q^{(0)}+Q^{(1)}$, mNGs are exact low-energy quasi-particles and thermal fluctuation broadens the peak in a symmetric way (blue). The inclusion of $Q^{(2)}$ causes a peak shift and also the lineshape becomes asymmetric (green). For Floquet systems without Trotter decomposition (purple), the center of the Fourier peak shifts again and starts to exhibit interesting fine structures, e.g., shoulder on the right side of the main peak. Although not shown, for a smaller driving period side peaks also arise. Hence, one can conclude that $Q^{(2)}$, that breaks all symmetries, is sufficient to generate asymmetric lineshape, but the fine structures, e.g. the splitting of the main peak, shoulders or side peaks, originate from $Q^{(3)}$ of order $\mathcal{O}(T^{3})$. This also qualitatively explains the scaling exponent in the splitting observed in Fig. S1(a). The 3rd order Trotter approximation leads to almost identical results (orange). Interestingly, the 2nd order Trotter decomposition leads to more visible side peaks.
@@
It is worth pointing out that, although the 2nd Trotterization being used in the main text results in a slightly different lineshape, compared the Runge-Kutta results, the overall linewidths are approximately the same, as shown in Fig. S2.
@@
Figure S2: The linewidth of the Fourier spectrum for different simulation methods. The Trotter protocol generates a slightly larger linewidth, but it closely mimics Runge-Kutta solver. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $J_{x}{=}1$ $\theta_{j}{=}0.25\pi$ and $\phi_{j}$ randomly sampled within $(-\pi/2,\pi/2)$. $100$ ensemble averages are performed to the system with its size $N=1000$.
4 Different Scaling of Linewidth

### v4

@@ Abstract
SM 1 Alternative Drive Protocols
1 Discrete Driving Protocol
Table 1: Different Protocols for Floquet-engineered mNGs
Protocol
Absence of Symmetry
in Instantaneous Hamiltonians
Hierarchical Symmetry Structure
in Effective Hamiltonian
Generator Protection in $\mathcal{O}(T)$ Driving Steps
[…]
ii $\times$ $\times$ -- 4
iii $\times$ 4
iv 8
In the main text, we propose a protocol that gives rise to mNGs with tunable mass and lifetimes in driven many-body systems. However, these features arise also in other time-dependent protocols. In this section, we discuss four versatile protocols with respective advantages and shortcomings concerning the experimental accessibility and the stability of mNGs, as summarized in Table 1. More precisely, we focus on the following features:
  • (1)
    No requirement of perfect symmetry in the instantaneous Hamiltonian generators: Modern quantum simulators do not realize perfectly symmetric Hamiltonians. Therefore, in practice, designing a Floquet scheme that does not require the implementation of perfect symmetries, is of great importance to Floquet engineering mNGs. Furthermore, at the fundamental level, if the instantaneous Hamiltonians do not preserve any symmetry, they should not host any mNGs. Therefore, the appearance of Floquet-engineered mNGs after the entire driving cycle, becomes a truly emergent non-equilibrium phenomenon, which does not exist without the drive.
  • (2)
    Hierarchical symmetry structure in effective Hamiltonian: This implies that, $Q^{(0)}$ preserves a non-abelian symmetry group $G_{2}$, $Q^{(1)}$ preserves a sub-group, $G_{1}\subset G_{2}$, and higher-order terms break all remaining symmetries. Even though the entire Floquet operator does not preserve any symmetry, this mechanism ensures the existence of mNGs in the high-frequency limit, and its mass is generated by $Q^{(1)}$ at the order $\mathcal{O}(T)$. However, the determination of the mass of mNGs can be complicated and strongly depends on the details of the microscopic Hamiltonian, see discussions in 8.
  • (3)
    Generator protection of the mass of mNGs in the effective Hamiltonian at the order $\mathcal{O}(T)$: We can go beyond the aforementioned hierarchical symmetry breaking construction and notice that, if $Q^{(1)}$ precisely corresponds to one of the symmetry generators of $G_{2}$, mNGs’ mass can be determined analytically by exploiting the non-abelian group structure. This feature is particularly important when the usual numerical analysis of mNGs faces difficulties, especially in the thermodynamic limit. It also allows for the exact controllability of the mass magnitude by the Floquet drive.
  • (4)
    Number of driving steps: The length of a driving protocol characterizes its complexity and short protocols are generally more accessible in experiments. For instance, in gate-based digital quantum platforms, more steps imply that more quantum gates are involved, which inevitably introduces extra gate errors that further destabilize the mNGs.
The detailed properties of the four protocols can be summarized as follows:
Protocol i. This is the protocol from Eq. (1) in the main text. This 14-step piece-wise protocol requires symmetry in the instantaneous Hamiltonians. Despite its complexity, its primary advantage lies in the structure of the effective Hamiltonian, where $Q^{(1)}$ corresponds to the symmetry generator of $G_{2}$. For sufficiently low-temperature initial states, the damping process of mNGs arises from higher-order contributions to the effective Hamiltonian $Q^{(2)}$. Additionally, the EOM generated by the instantaneous Hamiltonian in this protocol is either analytically integrable or can be conveniently decomposed into integrable parts using Trotter decomposition, which significantly facilitates the efficiency of numerical simulations, see details in SM 5.
Protocol ii. One can try to simplify the protocol above by considering the 4-step drive
$U_{F}=U_{0}^{+}U_{2}U_{M}U_{0}^{-}=e^{-iTH_{0}}e^{-iH_{2}T}e^{-iH_{M}\theta T}e^{iH_{0}T},$ (S.1)
where $H_{2}$ preserves the $G_{2}$ symmetry, $H_{M}$ is a generator of $G_{2}$ and preserves $G_{1}\subset G_{2}$, and $H_{0}$ breaks all symmetries. $U_{M}$ is a global rotation around the generator $H_{M}$ for a time-window $\theta T$. Thus, the leading order of the effective Hamiltonian reads
$Q^{(0)}=H_{2}+\theta H_{M},$ (S.2)
and mNGs exist with the mass linearly depending on $\theta$. However, unlike Protocol i, it does not allow for a hierarchical symmetry breaking process, and the symmetry $G_{1}$ is already broken by $Q^{(1)}$ of order $\mathcal{O}(T)$.
Both Protocol i and ii require the direct access of the Hamiltonian $H_{2}$ that preserves a non-abelian symmetry $G_{2}$. This can be particularly challenging for state-of-the-art quantum simulator platforms. In practice, the accessible instantaneous Hamiltonian can take the following form, $H=H_{n}+H_{n-1}+\cdots+H_{0}$, where $H_{0}$ breaks all symmetries. Yet, flipping some of the signs in front of each Hamiltonian can be easily achieved, for instance, by applying single-site $\pi/2$ gates on Rydberg atom quantum simulators [50]. Next, we introduce two Floquet protocols by combining single-site gates and the many-body Hamiltonian $H$, such that Floquet-engineering mNGs become experimentally feasible.
Protocol iii. We consider the 4-step drive
$U_{F}=U_{(+1,+1)}U_{(-1,-1)}U_{(+1,-1)}U_{(-1,+1)}$ (S.3)
where $U_{(\tau_{1},\tau_{2})}=e^{-i(H_{2}+\tau_{1}\cdot H_{1}+\tau_{2}\cdot H_{0})T},\ \tau_{i}=\pm 1$, $H_{2}$ preserves $G_{2}$ symmetry, $H_{1}$ preserves $G_{1}\subset G_{2}$, and $H_{0}$ breaks all symmetries. Its effective Hamiltonian reads $Q^{(0)}\propto H_{2}$, $Q^{(1)}\propto iT[H_{2},H_{1}]$, which features a hierarchical symmetry structure, and $G_{1}$ is broken by $Q^{(2)}$. However, $Q^{(1)}$ does not correspond to the generator of $G_{2}$ and, therefore, the prediction of the mNG mass can be difficult. We resolve this issue by considering the following protocol.
Protocol iv. The protocol reads
$U_{F}=$ $U_{(+1,+1,+1)}U_{(+1,-1,-1)}U_{(-1,-1,-1)}U_{(-1,+1,+1)}$ (S.4)
$U_{(-1,-1,+1)}U_{(-1,+1,-1)}U_{(+1,+1,-1)}U_{(+1,-1,+1)}$
where $U_{(\tau_{1},\tau_{2},\tau_{3})}=e^{-i(H_{2}+\tau_{1}\cdot H_{1}+\tau_{2}\cdot H_{1}^{\prime}+\tau_{3}\cdot H_{0})T},\ \tau_{i}=\pm 1$, and the Hamiltonians $H_{2}$, $H_{1}$, $H_{1}^{\prime}$ and $H_{0}$ are the same as Eq. (1). Its effective Hamiltonian reads, $Q^{(0)}\propto H_{2}$, $Q^{(1)}\propto iT[H_{1},H_{1}^{\prime}]$, and $Q^{(1)}$ is a generator of $G_{2}$. This protocol combines the advantages of both protocol i and protocol iii, allowing for an exact prediction of mass, while still being experimentally feasible.
In summary, we now provide four different protocols and discuss their advantages and disadvantages, in terms of both their experimental feasibility and the properties of the realized mNGs. Importantly, we propose protocol iii and protocol iv  that do not require direct access to non-abelian many-body Hamiltonians, demonstrating the feasibility of Floquet engineered mNGs for state-of-the-art quantum simulator platforms. This variety of different protocols also highlights that our proposed Floquet-engineered mNGs is a fundamental and fairly universal emergent phenomenon.
2 Continuous Driving Protocol
@@
$Q^{(0)}\propto K_{0},\qquad Q^{(1)}\propto T\sum_{l=1}^{\infty}\frac{1}{l}[K_{l},K_{-l}],$ (S.6)
@@
where $K_{0}$ preserves a symmetry group $G_{2}$, $K_{1}$ and $K_{1}^{\prime}$ are two non-commuting generators of $G_{2}$, and $K_{2}$ further breaks all symmetries. Then the effective Hamiltonian reduces to $Q^{(0)}\propto K_{0},\ Q^{(1)}\propto iT[K_{1},K_{1}^{\prime}]$ and higher-order terms break all symmetries. Hence, this protocol can effectively realize mNGs similar to those in the main text but with continuous drives.
@@
$\dot{S}_{j}^{z}=$ 0 0\ .0 .
This can be solved analytically by performing a Fourier transform, $S_{j}^{\alpha}{=}\sum_{q}S_{q}^{\alpha}e^{i(qa)j}$, $\alpha{\in}\{x,y,z\}$, with $q$ denoting the quasi-momentum and $a$ being the lattice constant. Finally, we get the spin-wave spectrum, Eq. (5)
@@ 2 Oscillating Amplitudes of Quasi-conserved Quantities in Classical Spin Systems
[…]
$H=-J_{2}\sum_{j}\bm{S}_{j}\cdot\bm{S}_{j+1}-h\left(h_{x}\sum_{j}S_{j}^{x}+h_{y}\sum_{j}S_{j}^{y}+h_{z}\sum_{j}S_{j}^{z}\right),$ (S.11)
@@
$\dot{\bm{S}}_{q}=[2J_{2}(1-\cos{qa})+h]\begin{pmatrix}0&h_{z}&-h_{y}\\ -h_{z}&0&h_{x}\\ h_{y}&-h_{x}&0\\ \end{pmatrix}\bm{S}_{q}\equiv A_{q}\bm{S}_{q}.$ (S.13)
@@
[…]
3 Nonlinear Effect
Zooming into Fig. 2, there exists a very faint line in DSF, which goes beyond the prediction of mNGs obtained by the linearized EOM. We attribute it to the nonlinear effects induced by the Heisenberg model, and the drive-induced higher-order corrections.
To show this, we first compare the DSF for the quenched dynamics of $Q_{\text{eff}}$ and the Floquet protocol. Fig. S1(a) depicts the contribution of the $x$ and $y$ components to the DSF, where only the gapped mode is visible. A gapless signature appears in Fig. S1(b), where the DSF of the $z$ component is plotted. The spin-wave solution of the linearized EOM, Eq. (S.9), predicts that the $z$-component should remain static. However, when the system is initialized with randomness, spatial fluctuations in z direction indeed become dynamical. Due to total magnetization conservation in the $z$-direction, the excitation spectrum is gapless. We now analytically solve for this spectrum by using a perturbative expansion w.r.t. the linearized solution. As shown on the right side of Fig. S1(b), it matches well with the most dominant signal obtained by numerical simulations.
Figure S1: Dynamical structure factor (DSF) over the Brillouin zone. (a) and (b) are obtained from spin-dynamics simulation of a time-independent Hamiltonian, Eq. (4). (a) contains the contribution of x and y components, and the gapped mNG mode with quadratic dispersion appears, matching well the theoretical prediction. (b) illustrates the DSF of the $z$-component, where a gapless signal is visible due to the nonlinear effect of the Heisenberg model, which also aligns with our analytical calculation. (c) corresponds to the DSF for the $z$-component obtained by the Floquet protocol, Eq. (1). Both the gapped mode and the gapless signal can be observed, since magnetizations in three directions are coupled via higher order drive-induced processes.
The mean-field solution, i.e., the linearized solution for $q=0$, reads
$\bar{S}^{x}(t)$ $=\bar{S}^{x}(0)\cos{(\omega_{\text{gap}}t)}+\bar{S}^{y}(0)\sin{(\omega_{\text{gap}}t)},$ (S.16)
$\bar{S}^{y}(t)$ $=\bar{S}^{y}(0)\cos{(\omega_{\text{gap}}t)}-\bar{S}^{x}(0)\sin{(\omega_{\text{gap}}t)},$
$\bar{S}^{z}(t)$ $=\bar{S}^{z}(0)$
where $\omega_{\text{gap}}=|h_{x}h_{y}|T/200$. We consider small fluctuations around the mean-field solution as $S_{j}^{\alpha}=\bar{S}^{\alpha}+\delta S_{j}^{\alpha}$ with $\delta S_{j}^{\alpha}\ll 1$. By expanding Eq. (S.8) to the leading order of $\delta S_{j}^{\alpha}$ , we obtain
$\delta\dot{S}_{j}^{x}=$ $\frac{J_{2}}{5}(2\delta S_{j}^{y}-\delta S_{j-1}^{y}-\delta S_{j+1}^{y})-\frac{h_{x}h_{y}T}{200}\delta S_{j}^{y}-\frac{J_{2}}{5}\bar{S}_{y}(2\delta S_{j}^{z}-\delta S_{j-1}^{z}-\delta S_{j+1}^{z}),$ (S.17)
$\delta\dot{S}_{j}^{y}=$ $\frac{J_{2}}{5}(\delta S_{j-1}^{x}+\delta S_{j+1}^{x}-2\delta S_{j}^{x})+\frac{h_{x}h_{y}T}{200}\delta S_{j}^{x}+\frac{J_{2}}{5}\bar{S}_{x}(2\delta S_{j}^{z}-\delta S_{j-1}^{z}-\delta S_{j+1}^{z}),$
$\delta\dot{S}_{j}^{z}=$ $\frac{J_{2}}{5}\bar{S}_{y}(2\delta S_{j}^{x}-\delta S_{j-1}^{x}-\delta S_{j+1}^{x})-\frac{J_{2}}{5}\bar{S}_{x}(2\delta S_{j}^{y}-\delta S_{j-1}^{y}-\delta S_{j+1}^{y}).$
Since the mean-field solution is valid for $\bar{S}^{x/y}\ll 1$, the last terms on the right-hand side of the first two rows can be neglected. Then, by performing a Fourier transform, $S_{j}^{\alpha}{=}\sum_{q}S_{q}^{\alpha}e^{i(qa)j}$, we obtain $\delta S_{q}^{x/y}=A_{q}^{x/y}e^{-i\omega t}$, where $\omega$ is the same as Eq. (S.10) and $A_{q}^{x}=iA_{q}^{y}$. Plugging the solution into the EOM of $\delta S_{q}^{z}$, we have $\delta\dot{S}_{q}^{z}\propto e^{i(\omega-\omega_{\text{gap}})t}$. Therefore, the spectrum of the $z$-components is quadratic and gapless
$\omega_{z}=\omega-\omega_{\text{gap}}=\pm\frac{2J_{2}}{5}(1-\cos{qa}).$ (S.18)
Furthermore, drive-induced higher-order effects slightly modify the spectrum. Crucially, as is mentioned in the main text, the explicit SB generator of the $O(2)$ symmetry slightly deviates from the $z$-direction due to higher-order corrections, which couple the fluctuations in all three directions. Therefore, in Floquet systems, DSF of the $z$-component (Fig. S1(c)) exhibits both a gapped and a gapless spectrum. Similarly, the DSF of the $x$ and $y$ components also exhibits a faint gapless spectrum in Fig. 2.
@@ 4 Asymmetric Lineshape and Side Peaks
The DSF has an asymmetric lineshape, as shown in Fig. 4(a). Also, several side peaks can be clearly resolved – an intriguing and rare non-equilibrium phenomenon in systems without external driving. We conjecture that the side peaks originate from $Q^{(3)}$ of order $\mathcal{O}(T^{3})$, a high-order and weak process. Their appearance can be slightly enhanced by the Trotter decomposition being used to approximate the Heisenberg Hamiltonian evolution. To demonstrate this, we compare the numerical simulation generated by $Q_{\text{eff}}$ and different order Trotter decompositions.
Figure S2: (a) The splitting between the main peak and the highest side peak $\Delta\omega$ exhibits a power-law scaling as $\Delta\omega\sim T^{\alpha}$, $\alpha\approx 3$, suggesting that these peaks may be generated by $Q^{(3)}$ of order $\mathcal{O}(T^{3})$. (b) Fourier peaks obtained by various numerical methods. We use $T{=}1.2$ and $J_{x}{=}1$. The lineshape becomes asymmetric once $Q^{(2)}$ is included. Fine structures in the lineshape appear in Floquet systems. 3rd-order Trotterization leads to a Fourier spectrum that is almost identical to those obtained without the Trotterization. 2nd order Trotter slightly enhances the side peaks, although the overall linewidth remains approximately unchanged. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $\theta_{j}{=}0.25\pi$ and $\phi_{j}$ randomly sampled within $(-\pi/2,\pi/2)$. 100100100 ensemble averages are performed for a system size of $N=1000$.
We extract the splitting $\Delta\omega$ between the main peak and the highest side peak and plot it in Fig. S2(a) for different driving periods. $\Delta\omega$ exhibits a power-law dependence on the driving period $T$ with the exponent $\alpha\approx 3$.
We then try to reproduce the interesting lineshape by several different numerical methods:
@@
2. Floquet driving the system and approximating the Heisenberg Hamiltonian $H_{2}$ using Trotter decomposition of different orders. Details of this method will be explained later in Sec. SM 5.
@@
The results are shown in Fig. S2(b). For $Q^{(0)}+Q^{(1)}$, mNGs are exact low-energy quasi-particles and thermal fluctuation broadens the peak in a symmetric way (blue). The inclusion of $Q^{(2)}$ causes a peak shift; moreover, the lineshape becomes asymmetric (green). For Floquet systems without Trotter decomposition (purple), the center of the Fourier peak shifts again and starts to exhibit interesting fine structures, e.g., the shoulder on the right side of the main peak. Although not shown, for a smaller drive period side peaks also arise. Hence, one can conclude that $Q^{(2)}$, which breaks all symmetries, is sufficient to generate an asymmetric lineshape, but the fine structures, e.g., the splitting of the main peak, shoulders or side peaks, originate from $Q^{(3)}$ of order $\mathcal{O}(T^{3})$. This also qualitatively explains the scaling exponent in the splitting observed in Fig. S2(a). The 3rd order Trotter approximation leads to almost identical results (orange). Interestingly, the 2nd order Trotter decomposition leads to more visible side peaks.
@@
It is worth pointing out that, although the 2nd Trotterization used in the main text results in a slightly different lineshape, compared to the Runge-Kutta results, the overall linewidths are approximately the same, as shown in Fig. S3.
@@
Figure S3: Linewidth of the Fourier spectrum for different simulation methods. The Trotter protocol generates a slightly larger linewidth, but it closely mimics the Runge-Kutta solver. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $J_{x}{=}1$ $\theta_{j}{=}0.25\pi$ and $\phi_{j}$ randomly sampled within $(-\pi/2,\pi/2)$. 100100100 ensemble averages are performed for a system size of $N=1000$.
5 Lifetime of Modes with Finite Momenta
In the main text, we investigate the lifetime of the $q=0$ mode. In Fig. 4, we find that the linewidth of the DSF scales as $T^{3}$, and we give one possible explanation using a hydrodynamic theory. Here, we further discuss the lifetime of modes with finite momenta.
For simplicity, we first consider the static system governed by the effective Hamiltonian $Q^{(0)}+Q^{(1)}$. In the case of the Heisenberg model with an additional z-field, different $q$ modes have infinite lifetime. Quantum mechanically, this is true since all single-magnon excitations are exact eigenstates of $Q^{(0)}+Q^{(1)}$. Classically, the spin-wave solution of the approximate linearized EOM is in fact an exact solution of non-linear EOM for the classical Heisenberg model. More generally, for the case in the main text of our work and also Ref. [9], where $Q^{(0)}$ preserves a non-Abelian symmetry $G_{2}$ and $Q^{(1)}$ is proportional to a generator of $G_{2}$, the infinite lifetime of modes with arbitrary momentum $q$ is also guaranteed. According to the Goldstone Theorem, there exists a gapless excitation for $Q^{(0)}$, $Q^{(0)}|q\rangle=E(q)|q\rangle$, where $E(q)$ denotes the eigenenergy for the single-magnon excitation at a fixed $q$. Due to $[Q^{(0)},Q^{(1)}]=0$, $Q^{(0)}$ and $Q^{(1)}$ share the same set of eigenstates, which implies that $Q^{(1)}|q\rangle=K(q)|q\rangle$. Therefore, mNGs are also infinitely long-lived with eigenenergy $E(q)+K(q)$.
This becomes more complicated if the symmetry-breaking term $Q^{(1)}$ is not a generator of $G_{2}$, hence $[Q^{(0)},Q^{(1)}]$ no longer vanishes. Therefore, $Q^{(0)}$ and $Q^{(1)}$ do not have simultaneous eigenstates. Construction of the mNG mode, e.g. by using spin-wave approximation, then depends on the specific form of the Hamiltonian. Their lifetime may also depend on $q$ and this certainly goes beyond the scope of this work.
Figure S4: The width for different $q$ modes. The linewidth of the $q=\pi/100$ mode (orange) exhibits a similar $T^{3}$ scaling behavior in the high-frequency limit as the $q=0$ mode. The black dashed line with scaling $T^{3}$ is a guide to the eye. The $q=\pi/5$ mode exhibits a broader linewidth, and hence a shorter lifetime. Its lifetime is largely independent of $T$, suggesting that its decay is dominated by $Q^{(0)}$ and the thermal fluctuations in the initial state. We use $J_{2}{=}1$, $h_{x}{=}2$, $h_{y}{=}{-}2$, $J_{x}{=}1$ $\theta_{j}{=}0.25\pi$ for numerical simulations and $\phi_{j}$ is randomly sampled within $(-\pi/2,\pi/2)$. 100100100 ensemble averages are performed for a system size of $N=1000$.
However, in our work, we always start from an initial state with spatial fluctuations at low but finite temperature. Floquet driving also induces higher-order effects that further destabilize the mNGs. Therefore, we numerically examine the dependence of the linewidth on momentum $q$ of the mode in Fig. S4. We use the same parameters and initial states as the yellow dots in Fig. 4(b) and extract the linewidth of two extra modes: one has a momentum close to zero while the other has a large momentum. The linewidth indeed depends on $q$ in a non-trivial way. For a fixed driving period $T$, larger $q$ mode exhibits a large linewidth, hence shorter lifetime. The $q=\pi/100$ mode shows a $T^{3}$ scaling, which is consistent with the hydrodynamics prediction that works particularly well in the long-wavelength limit. For $q=\pi/5$ mode, the linewidth is largely independent of $T$, indicating that mNGs are destabilized by thermal fluctuations in the initial state and $Q^{(0)}$ that is independent of $T$, rather than the driving-induced perturbations. Intuitively, one can understand this behavior by noting that the Floquet protocol only involves local fields or interactions, hence it takes longer time to destroy the long range order, and excitations at short lengthscales are less stable. A deeper understanding of the heating behavior for different $q$ modes would require further investigation, which we leave for future work.
6 Different Scaling Exponent of Linewidth